OpenStudy (anonymous):

tan(x)-3cot(x)=0 (sin(x)/cos(x))-3(cos(x)/sin(x)) (sin^2(x)/(cos(x))(sin(x))-3(cos^2(x)/(cos(x))(sin(x)) so where do I go from here?

OpenStudy (anonymous):

\[\begin{align*}\tan x-3\cot x&=0\\ \frac{\sin x}{\cos x}-3\frac{\cos x}{\sin x}&=0\\ \cos x\left(\frac{\sin x}{\cos x}-\frac{3\cos x}{\sin x}\right)&=0\cos x\\ \sin x-\frac{3\cos^2 x}{\sin x}&=0\\ \sin x\left(\sin x-\frac{3\cos^2x}{\sin x}\right)&=0\sin x\\ \sin^2x-3\cos^2x=0 \end{align*}\] Next, use the identity \(\sin^2x+\cos^2x=1\): \[(1-\cos^2x)-3\cos^2x=0\\ 1-4\cos^2x=0\] which you can solve by factoring: \[(1-2\cos x)(1+2\cos x)=0\] or by some simpler rewriting: \[\cos^2x=\frac{1}{4}\\ \cos x=\pm\frac{1}{2}\]